OpenStudy (yato_gami):

I need help with this please! I want someone to go through the steps with me, not tell me the answer. Thanks~! ^_^ Which of the following is an equivalent function to the function below? f(x) = 30(1.03)3x

OpenStudy (yato_gami):

A. f(x) = 30(3.0900)x B. f(x) = 30(1.0927)x C. f(x) = 10(1.0927)x D. f(x) = 10(3.0900)x

OpenStudy (tsukiyama_shuu):

YATO

OpenStudy (yato_gami):

Tsukiyama!

OpenStudy (tsukiyama_shuu):

B. f(x) = 30(1.0927)^x

OpenStudy (yato_gami):

Ok cool! Can you tell me how ya got it? I wanna understand how to do it. ^_^

OpenStudy (tsukiyama_shuu):

http://brainly.com/question/1566588

OpenStudy (tsukiyama_shuu):

I cheated....

OpenStudy (yato_gami):

lolol welp! Thank anyway! ^_^

OpenStudy (tsukiyama_shuu):

Np ^^